# Three-particle scattering amplitudes from lattice QCD

### v1

@@ 1 Introduction
One question that the non-expert may well have is this: Why three? What about processes involving more than three particles? After all, many resonances have decay channels involving four or more particles. I am restricting the discussion to three particles simply because the necessary general formalism for going beyond three does not yet exist.
@@ 2 Status of formalism
The situation facing a LQCD practitioner desirous of calculating scattering amplitudes is this: LQCD is very effective at determining the spectrum of finite-volume states in spatial boxes of size 2102-10\,fm, using Euclidean correlation functions, but the scattering amplitudes that we want are infinite-volume quantities involving in and out states. How are the spectrum and amplitudes related? In this regard, it is important to keep in mind that the finite-volume energies are physical quantities (after taking the continuum limit), so that we are asking for a relation between one set of physical quantities and another. And it is plausible that, for a box large enough that the hadrons do not significantly overlap, the finite-volume energies depend on the scattering amplitudes. The question is how to extract this information.
Lüscher solved this for the case of two particles [13, 14, 15]. This works as follows (using a notation convenient for the generalization to three particles). Given the two-particle K matrix, $\mathcal{K}_{2}$, an infinite-volume quantity, the finite-volume spectrum is given by the values of $E$ that solve the QC2,
@@
Here $\boldsymbol{P}$ is the (quantized) total momentum in the box, which I assume is cubic with length LL. [^3] The generalization to other spatial box shapes is straightforward [24, 25]. Thus the center of momentum (CM) energy is given by $E^{*}=\sqrt{E^{2}-\boldsymbol{P}^{2}}$, and it is this that the two-particle K matrix depends on. Both quantities in eq.˜1 are matrices in $\{\ell,m\}$ space, where $\ell$ is the relative angular momentum of the two particles in the CM frame, and $m$ the azimuthal quantum number. $\mathcal{K}_{2}$ is diagonal in $\ell$ and $m$, whereas the purely kinematical function $F_{2}$ (the “Lüscher zeta function”) is, in general, not diagonal, reflecting the violation of rotational symmetry by the finite box. [^4] Explicit expressions for $F_{2}$, and the relation of $\mathcal{K}_{2}$ to phase shifts, are given, e.g., in Ref. [26].
@@
A potentially more important source of error comes from the fact that, to use the QC2, one must truncate the angular momentum index $\ell$ in order to have finite matrices. This is achieved formally by setting $\mathcal{K}_{2}$ to zero for $\ell>\ell_{\rm max}$. Near to threshold, barrier factors lead to $\mathcal{K}_{2}\propto q^{2\ell}$, so such a truncation is reasonable. At higher energies it is a less controlled approximation, though it is one that is commonly used also in analysis of experimental data.
@@
where $\rho$ is the phase-space matrix (given explicitly in, e.g., Ref. [26]; one should beware, however, that different definitions of K matrices are used in literature). Given that $\mathcal{K}_{2}$ is an hermitian matrix, this relation ensures that $\mathcal{M}_{2}$ is unitary, and, in particular, includes the phases induced by initial- and final-state interactions. Using this result, and the parametrization of $\mathcal{K}_{2}$, one can then analytically continue onto the second Riemann sheet (or sheets, in the case of multiple two-particle channels), and determine resonance parameters.
@@
I now turn to the three-particle system. The formalism was developed first for three identical scalar particles with a G-parity-like symmetry, following three different approaches. The first used an all orders diagrammatic analysis in a relativistic effective field theory (EFT) describing generic interactions between the particles, and is dubbed the RFT approach [28, 29]. This approach makes no assumption about the form of the interactions, nor does it rely on a power counting. The second used nonrelativistic (NR) EFT, and led to a much simpler derivation [30, 31], although the three-particle interaction was initially restricted to the leading order terms in the NR power counting. Subsequent work has shown how to relativize this approach [32]. The third approach used a unitary representation of the three-particle amplitude developed in the continuum [33], and from that developed a finite-volume quantization condition [34]. It is denoted the “finite-volume unitarity” or FVU approach.
@@
The workflow for determining the three-particle scattering amplitude consists, in all approaches, of two steps. I will describe these for the case in which there are no transitions between two and three particles, e.g. three pions in isosymmetric QCD. In the first step, one fits the finite-volume energies for two and three-particles (which, by assumption, can be calculated separately) to the predictions of the two- and three-particle quantization conditions, respectively (QC2 and QC3), and thus determines (the parametrized forms) of the two- and three-particle K matrices. In the second step, one solves integral equations in which the K matrices are kernels, the output of which is the three-particle scattering amplitude, $\mathcal{M}_{3}$. The QC3 has the form
@@
As for the QC2, solutions to the QC3 give the values of $E$ at which there are finite-volume energies for the given K matrices. All quantities in these equations are matrices in a space with indices $\{k,\ell,m,x\}$—a much larger space than in the QC2.666I stress that all three approaches to the three-particle formalism use matrices with the same index space. The formalism separates the three particles into a spectator, with momentum $\boldsymbol{k}$, and the remaining pair. The index $k$ is a shorthand for the finite-volume momentum of the spectator, $\boldsymbol{k}\in(2\pi/L)\mathbb{Z}^{3}$. The indices $\ell,m$ represent the relative angular momentum of the particles in the pair in their CM frame. Finally, the mystery index $x$ depends on the system, and includes the effects of flavor and spin. It always takes a finite number of values, and is absent in the case of three identical particles.
@@
$\mathcal{M}_{3}=\mathcal{D}+\mathcal{M}_{\rm df,3}\,$, (5)
@@ 3.5 Dreaming of three neutrons
In fig.˜7, I show what happens to the 13 $G_{1}$ levels in the upper “band” from fig.˜6 as the coefficient of one of the terms in $\mathcal{K}_{\mathrm{df},3}$}, $c_{A}$, is varied from large negative to large positive values. The levels shift down by one step, as indicated by the vertical dotted lines. There is much more that can be said about these plots (see Ref. [115]), but the key message here is that the impact of $\mathcal{K}_{\mathrm{df},3}$} can be as large as the splittings induced by two-particle interactions. Thus a precision of $\sim 5\;$MeV will allow one to place significant constraints on three-neutron interactions.
@@ 4 Summary and Outlook
  • Moving to four or more particles. Doing so is clearly necessary for many applications, e.g. studying $\pi(1300)$ for physical masses, a complete treatment of the $\rho$ (since the $4\pi$ decay is kinematically allowed), and a method for calculating weak decays of $D$ mesons (where decays to two, three, four and more particles are allowed). One question here is whether the QC4, QC5, …path is feasible. Aside from the threshold expansion, the four-particle formalism is not yet known, although some pioneering efforts were described in the talk by Muhkerjee. I can envisage two possibilities. The optimistic one is that the impact of short-ranged interactions of four or more particles is very small, and a set of recursive quantization conditions and integral equations involving only $\mathcal{K}_{2}$ and $\mathcal{K}_{\mathrm{df},3}$} can be derived and made practical. The pessimistic one is that the QC4 turns out to be barely practical, while QCN (with $N>4$) is out of reach.
@@
There is, indeed, much work to be done, but, for me, this is a very worthwhile and, indeed, exciting, endeavor, with the potential to greatly improve our understanding of the strong interactions.

### v2

@@ 1 Introduction
One question that the non-expert may well have is this: Why three? What about processes involving more than three particles? After all, many resonances have decay channels involving four or more particles. There is, indeed, much interest in extending the work to more than three particles (a point that I return to at the end of this talk), but the results available to date have concerned three particles. It has turned out that the three-particle sector has been richer and more complicated than I, at least, expected.
@@ 2 Status of formalism
The situation facing a LQCD theorist desirous of calculating scattering amplitudes is this: LQCD is very effective at determining the spectrum of finite-volume states in spatial boxes of size 2102-10\,fm, using Euclidean correlation functions, but the scattering amplitudes that we want are infinite-volume quantities involving in and out states. How are the spectrum and amplitudes related? In this regard, it is important to keep in mind that the finite-volume energies are physical quantities (after taking the continuum limit), so that we are asking for a relation between one set of physical quantities and another. And it is plausible that, for a box large enough that the hadrons do not significantly overlap, the finite-volume energies depend on the scattering amplitudes. The question is how to extract this information.
Lüscher solved this for the case of two particles [13, 14, 15]. This works as follows (using a notation convenient for the generalization to three particles). Given the two-particle K matrix, $\mathcal{K}_{2}$, an infinite-volume quantity, the finite-volume spectrum is given by the values of $E$ that solve the QC2
@@
Here $\boldsymbol{P}$ is the (quantized) total momentum in the box, which I assume is cubic with length LL. [^3] The generalization to other spatial box shapes is straightforward [24, 25]. Thus the center of momentum (CM) energy is given by $E^{*}=\sqrt{E^{2}-\boldsymbol{P}^{2}}$, and it is this that the two-particle K matrix depends on. Both quantities in eq.˜1 are matrices in $\{\ell,m,x\}$ space, where $\ell$ is the relative angular momentum of the two particles in the CM frame, $m$ is the corresponding azimuthal quantum number, and the mystery index $x$ includes the effects of spin and multiple two-particle channels, and takes values that depend on the system [27]. $\mathcal{K}_{2}$ is diagonal in $\ell$ and $m$, whereas the purely kinematical function $F_{2}$ (the “Lüscher zeta function”) is, in general, not diagonal, reflecting the violation of rotational symmetry by the finite box. [^4] Explicit expressions for $F_{2}$, and the relation of $\mathcal{K}_{2}$ to phase shifts, are given, e.g., in Ref. [26].
@@
A potentially more important source of error comes from the fact that, to use the QC2, one must truncate the angular momentum index $\ell$ in order to have finite matrices. This is achieved formally by setting $\mathcal{K}_{2}$ to zero for $\ell>\ell_{\rm max}$. Near to threshold, barrier factors lead to $\mathcal{K}_{2}\propto q^{2\ell}$, where $q$ is the relative momentum of the two particles in their center of momentum frame, so such a truncation is reasonable. At higher energies it is a less controlled approximation, though it is one that is commonly used also in analysis of experimental data.
@@
where $\rho$ is the phase-space matrix (given explicitly in, e.g., Ref. [26]; one should beware, however, that different definitions of K matrices are used in literature). Given that $\mathcal{K}_{2}$ is an hermitian matrix, this relation ensures that $\mathcal{M}_{2}$ leads to a unitary S matrix, and, in particular, includes the phases induced by initial- and final-state interactions. Using this result, and the parametrization of $\mathcal{K}_{2}$, one can then analytically continue onto the second Riemann sheet (or sheets, in the case of multiple two-particle channels), and determine resonance parameters.
@@
I now turn to the three-particle system. The formalism was developed first for three identical scalar particles with a G-parity-like symmetry, following three different approaches. The first used an all orders diagrammatic analysis in a relativistic effective field theory (EFT) describing generic interactions between the particles, and is dubbed the RFT approach [28, 29]. This approach makes no assumption about the form of the interactions, nor does it rely on a power counting. The second used nonrelativistic (NR) EFT, and led to a much simpler derivation [30, 31], although the three-particle interaction was initially restricted to the leading order terms in the NR power counting. Subsequent work has shown how to relativize this approach [32]. The third approach used a unitary representation of the three-particle amplitude developed in the infinite volume [33], and from that developed a finite-volume quantization condition [34]. It is denoted the “finite-volume unitarity” or FVU approach.
@@
The workflow for determining the three-particle scattering amplitude consists, in all approaches, of two steps. I will describe these for the case in which there are no transitions between two and three particles, e.g. three pions in isosymmetric QCD. In the first step, one fits the finite-volume energies for two and three-particles (which, by assumption, can be calculated separately) to the predictions of the two- and three-particle quantization conditions, respectively (QC2 and QC3), and thus determines (the parametrized forms) of the two- and three-particle K matrices. In the second step, one solves integral equations in which the K matrices are kernels, the output of which is the three-particle scattering amplitude, $\mathcal{M}_{3}$.
The QC3 appearing in the first step has the form
@@
As for the QC2, solutions to the QC3 give the values of $E$ at which there are finite-volume energies for the given K matrices. All quantities in these equations are matrices in a space with indices $\{k,\ell,m,x\}$—a larger space than in the QC2.666I stress that all three approaches to the three-particle formalism use matrices with the same index space. The formalism separates the three particles into a spectator, with momentum $\boldsymbol{k}$, and the remaining pair. The index $k$ is a shorthand for the finite-volume momentum of the spectator, $\boldsymbol{k}\in(2\pi/L)\mathbb{Z}^{3}$. The indices $\ell,m$ represent the relative angular momentum of the particles in the pair in their CM frame. Finally, the index $x$ plays a similar role to that in the QC2, and accounts for the effects of flavor, spin, and the possibility of multiple three-particle channels. It always takes a finite number of values, and is absent in the case of three identical particles.
@@
$\mathcal{M}_{3}=\mathcal{D}(\mathcal{K}_{2})+\mathcal{M}_{\rm df,3}(\mathcal{K}_{2},{\mathcal{K}_{\mathrm{df},3}})\,$, (5)
@@ 3.5 Dreaming of three neutrons
In fig.˜7, I show what happens to the 13 $G_{1}$ levels in the upper (left-hand) “band” from fig.˜6 as the coefficient of one of the terms in $\mathcal{K}_{\mathrm{df},3}$}, $c_{A}$, is varied from large negative to large positive values. The levels shift down by one step, as indicated by the vertical dotted lines. There is much more that can be said about these plots (see Ref. [115]), but the key message here is that the impact of $\mathcal{K}_{\mathrm{df},3}$} can be as large as the splittings induced by two-particle interactions. Thus a precision of $\sim 5\;$MeV will allow one to place significant constraints on three-neutron interactions.
@@ 4 Summary and Outlook
  • Moving to four or more particles. Doing so is clearly necessary for many applications, e.g. studying $\pi(1300)$ for physical masses, a complete treatment of the $\rho$ (since the $4\pi$ decay is kinematically allowed), and a method for calculating weak decays of $D$ mesons (where decays to two, three, four and more particles are allowed). One question here is whether the QC4, QC5, …path is feasible. Aside from the threshold expansion, the four-particle formalism is not yet known, although some pioneering efforts were described in the talk by Mukherjee. I can envisage two possibilities. The optimistic one is that the impact of short-ranged interactions of four or more particles is very small, and a set of recursive quantization conditions and integral equations involving only $\mathcal{K}_{2}$ and $\mathcal{K}_{\mathrm{df},3}$} can be derived and made practical. The pessimistic one is that the QC4 turns out to be barely practical, while QCN (with $N>4$) is out of reach.
@@
There is clearly much work to be done, but, for me, this is a very worthwhile and exciting endeavor, with the potential to greatly improve our understanding of the strong interactions.